# When Should We Introduce Safety Interventions During Pretraining?

### v1

@@ Appendix A Additional Experiments
Additional Separability Results
@@ Appendix C Additional Discussion
One limitation of our experiments is that the introduction of safety pretraining interventions at different times during pretraining means that certain models have seen a larger number of tokens for synthetic contextualized data or refusal data. Thus, our results do not control the exact number of tokens from each pretraining mixture source. Finding an optimal “ordering” for a curriculum remains an interesting open question.
We also note that the notion of safety that we focus on in this paper is rather general, taken from the taxonomy proposed by the MLCommons [Vidgen et al., 2024]. This captures a wide variety of potential harms and undesirable behaviors. Our method likely requires additional efforts to tackle safety for specific capabilities, such as bio-weapons and cybersecurity [Li et al., 2024] or reliability of agents in context-dependent scenarios [Andriushchenko et al., 2025, Sam et al., 2025c].
@@ Appendix D Computational Resources
To perform pretraining for each of the 1.7B parameter models on 600B tokens, we used 4 nodes of 8xH100 GPUs for roughly 6-7 days. Finetuning requires roughly 4 H100 GPUs for less than 4 hours. Evaluations are run on a single NVIDIA L40S GPU in less than an hour.

### v2

@@ Appendix A Additional Experiments
Additional Separability Results.
@@
SafeBeam Overrefusal.
Figure 9 extends our overrefusal analysis to the SafeBeam inference setting. Compared to standard top-$k$ sampling, SafeBeam introduces a fifth response category (shortcircuit), which occurs when generation terminates upon producing the harmful metadata token. In deployment, such responses can be handled by substituting a predetermined safe completion.
@@ Appendix B Additional Experimental Details
Capabilities Evaluation
We use the default experimental setting for all the datasets as provided as in the LM evaluation harness [Gao et al., 2024]. We use the dataset default few-shot settings: 0-shot for ARC-Challenge, OpenBookQA, CommonsenseQA, PIQA, TriviaQA, and WinoGrande (num_fewshot=0), and 5-shot for GSM8K (num_fewshot=5; few-shot examples sampled from the training split).
@@ Appendix C Additional Discussion
We note that the notion of safety that we focus on in this paper is fairly broad, encompassing general notions of toxicity and harms. This is taken from the taxonomy proposed by the MLCommons [Vidgen et al., 2024], captures a wide variety of potential harms and undesirable behaviors in different domains. Our method likely requires additional efforts to tackle safety for specific capabilities, such as bio-weapons and cybersecurity [Li et al., 2024] or reliability of agents in context-dependent scenarios [Andriushchenko et al., 2025, Sam et al., 2025c].
@@ Appendix D Computational Resources
For each variant of the pretraining runs for the 1.7B parameter model on 600B tokens, we used 32 H100 GPUs for roughly 6-7 days. Finetuning requires roughly 4 H100 GPUs for 4 hours. Evaluations are run on one L40S GPU.